# Advances and Open Challenges in Federated Learning with Foundation Models

### v2

@@ VIII-D Promising Direction: Evaluation of FedFM
IX Conclusions
In this paper, we provide a comprehensive and perspective survey on the development of FedFM, representing a significant step towards building FM-based AI systems in an efficient and collaborative manner. The survey efforts reveal a complex yet promising set of topics intersecting FL and FMs. As FedFM continue to evolve, the focus on training and aggregation methodologies, trustworthiness measures, incentive mechanisms, and participant management strategies will be crucial for its success. This field is essential for unlocking the full potential of FL in leveraging the power of FMs. The proposed taxonomy for FedFM and novel strategies for participant selection and model training pave the way for future research, particularly in enhancing privacy-preserving techniques and leveraging quantum computing for improved security and efficiency. Overall, the convergence of FedFM demonstrates significant potential in advancing AI capabilities while effectively addressing critical challenges in data privacy and model trustworthiness.

### v4

@@ IX-D Promising Direction: Model Evaluation of FedFM
X Conclusions
In this paper, we provide a comprehensive and perspective survey on the development of FedFM, representing a significant step towards building FM-based AI systems in an efficient and collaborative manner. The survey reveals a complex yet promising set of topics intersecting FL and FMs. As FedFM continue to evolve, the focus on training and aggregation methodologies, trustworthiness measures, incentive mechanisms, and participant management strategies will be crucial for its success. This field is essential for unlocking the full potential of FMs by enabling efficient and trustworthy access to privately owned valuable data via FL. The proposed taxonomy for FedFM and novel strategies for participant selection and model training pave the way for future research, particularly in enhancing privacy-preserving techniques and leveraging quantum computing for improved security and efficiency. Overall, the convergence of FedFM demonstrates significant potential in advancing AI capabilities while effectively addressing critical challenges in data privacy and model trustworthiness.